• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gilman

    Log entry time 01:23:42 on March14,2012

    Entry number 366276

    keyword=shut valve

    MCC calls and asks about a target exit valve 1h04c drifting shut,
    They have trouble getting it open. We do not see such a valve on the target, but we notice that the left spectrometer n2 level CLL71Q310 had gone to 0. A few minutes later MCC calls and says they have managed to open it. See stripchart.

    FIGURE 1